# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 3.2 Bariant’s Attack
Recently Bariant introduced an attack on Ciminion via univariate greatest common divisor computations [Bar23]. As in our polynomial model he introduced an auxiliary variable $X$ for the truncated component. If one inverts $p_{E}$ and $p_{C}$ using $(c_{1},c_{2},X)$ as output state one then obtains three polynomials, see [Bar23, Fig. 3],
@@
Since the first polynomial is univariate in $X$ one can compute its greatest common divisor with $X^{q}-X$ to solve for $X$ and henceforth also solve for $K_{1}$ and $K_{2}$.
We note that this attack can also be interpreted in terms of term order conversion of our DRL Gröbner basis to LEX, since the univariate polynomial $\aleph=f_{1}(X)$ is simply the univariate polynomial of the Ciminion LEX Gröbner basis in $x$-shape position. Moreover, the construction of the univariate LEX polynomial can be effectively expressed in the cost of univariate polynomial multiplication over finite fields [HvdH22] which is quasi-linear in the degree.
As countermeasure Bariant proposes to add a linear combination of $K_{1}$ and $K_{2}$ to each component after $p_{E}$. We note that this modification breaks our Gröbner basis from LABEL:Th:_Ciminion_Gr\"obner_basis. As settlement, we propose to add a linear combination of $K_{1}$ and $K_{2}$ only to the non-linear parts of the round functions, e.g.
$\begin{split}\mathcal{R}^{(l)}:\mathbb{F}_{q}^{3}\times\mathbb{F}_{q}^{2}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix},(K_{1},K_{2})&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&0&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+\alpha\cdot K_{1}+\beta\cdot K_{2}\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},\end{split}$ (34)
where $\alpha,\beta\in\mathbb{F}_{q}$ and $\alpha\cdot\beta\neq 0$. This modification does not affect the proof of LABEL:Th:_Ciminion_Gr\"obner_basis, hence we maintain the DRL Gröbner basis for the modified Ciminion polynomial system. However, this modification breaks the construction of the univariate LEX polynomial via univariate polynomial multiplication, because the DRL Gröbner basis polynomial for the $r_{C}$th round will now also contain linear terms for the variables $x_{C,3}^{(2)},\dots,x_{C,3}^{(r_{C}-2)}$ instead of only $x_{C,3}^{(r_{C}-1)},x_{E,3}^{(1)},\dots,x_{E,3}^{(r_{E})},x$. While the LEX Gröbner basis of the modified Ciminion might still be in $x$-shape position, we now can only construct it by running a FGLM algorithm.
Note if the modification of Equation 34 is performed in all rounds of $p_{E}$, then the Middle State-Output Relation polynomial system of Bariant et al. [BBLP22, §5] becomes underdetermined, so this attack connecting two outputs of two nonce samples would not be applicable anymore.
@@ 3.3 Implementation
We implemented Ciminion in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^3] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In principle, the SageMath implementation can handle $128$ bit prime numbers, though we do not recommend using primes of that size for experimental Gröbner basis computations, because SageMath uses Python implementations for primes of that size. The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
We implemented the Gröbner basis computation from LABEL:Th:_Ciminion_Gr\"obner_basis in SageMath. For practical parameters $p=2^{127}+45$, $r_{C}=90$ and $r_{E}=14$ a DRL Gröbner basis can be computed in negligible time.
@@ 3.3.1 Ciminion2
We also implemented a modified version of Ciminion to protect against Bariant’s attack, see Equation 34, called Ciminion2, which adds $K_{1}+K_{2}$ in the last round of $p_{C}$ and every round of $p_{E}$ to the non-linear component. I.e., the round function modifies to
$\mathcal{R}^{(l)}:\begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(l)}&c_{4}^{(l)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+K_{1}+K_{2}\end{pmatrix}+\begin{pmatrix}c_{1}^{(l)}\\ c_{2}^{(l)}\\ c_{3}^{(l)}\end{pmatrix},$ (35)
@@
where $r_{C}\leq l\leq r_{C}+r_{E}$. For Ciminion2 a DRL Gröbner basis is also computed in negligible time in SageMath for practical parameters.
@@ 4 Analysis of Hydra
is affine equivalent to the Feistel permutation
@@
via the matrices $\mathbf{A}_{1}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ and $\mathbf{A}_{2}=\begin{pmatrix}1&1\\ 0&1\\ \end{pmatrix}$, i.e. $\mathcal{L}=\mathbf{A}_{2}\circ\mathcal{F}\circ\mathbf{A}_{1}$. Inspired by this equivalence one can reduce the Hydra polynomial system to only one non-linear equation per round, i.e. the Hydra polynomial system transforms to a Feistel-like polynomial system.
In [Ste24, §6.2] the iterated polynomial systems representing block ciphers based on generalized Feistel networks were analyzed. For these polynomial systems linear systems were derived, and in case the systems are of full rank the iterated polynomial systems are in generic coordinates, see [Ste24, Theorem 6.5].
Based on the aforementioned affine equivalence, we are going to replicate this strategy to construct such a linear system for Hydra.
@@ 4.1 Generic Coordinates for Hydra
Recall that the Hydra iterated polynomial system (Definition 2.2) consists of $816\cdot r_{\mathcal{H}}+$ 8 many polynomials in $416\cdot r_{\mathcal{H}}+$ 4 many variables. Let
$\mathbf{A}_{n}=\begin{pmatrix}1&0&\ldots&0&-1\\ 0&1&\ldots&0&-1\\ &&\ddots&&\\ 0&0&\ldots&1&-1\\ 0&0&\ldots&0&1\end{pmatrix}\in\operatorname{GL}_{n}\left(\mathbb{F}_{q}\right),$ (38)
and let $\mathbf{B}=\operatorname{diag}\left(\mathbf{A}_{4},\mathbf{A}_{4}\right)\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$. Then the first seven components of $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$, where $1\leq i\leq r_{\mathcal{H}}$ and $j\in\{1,2\}$, are affine and the eighth component is quadratic. Moreover, the forth and the eighth components of $\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}$ are quadratic while the other ones are affine. In particular, by extracting the $2\cdot 7\cdot r_{\mathcal{H}}$ affine components of the $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$’s we can eliminate the variables
@@
where $10\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$, in the quadratic equations. (This can be explicitly seen in the proof of Theorem 4.1.) Thus, after the elimination we reduce to $82\cdot r_{\mathcal{H}}+$ 8 polynomials, of whom $6$ are also affine, in $42\cdot r_{\mathcal{H}}+$ 4 many variables. We call this polynomial system the downsized Hydra polynomial system.
Analog to the analysis of Feistel networks in [Ste24, §6.2], we now derive a linear system for Hydra with which we can verify whether the Hydra polynomial system is in generic coordinates.
@@
$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{R}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
@@
Then every homogeneous ideal $I\subset P[x_{0}]$ such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\subset I$ is in generic coordinates if the following linear system has rank $416\cdot r_{\mathcal{H}}+$ 4
@@
$x_{j,k}^{(i)}-x_{j,8}^{(i)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i+1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i+1)}-\mathbf{k}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}$ $=0,$
$x_{j,k}^{(r_{\mathcal{H}}-1)}-x_{j,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
$x_{j,k}^{(r_{\mathcal{H}}-1)}-x_{j,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
$y_{1}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{1}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
@@
$y_{3}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{3}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}$ $=0,$
@@
By [Ste24, Theorem 3.2] we have to show that
$\sqrt{\mathcal{G}^{\text{top}}}=\left(\mathbf{y},\mathbf{z},\mathbf{x}_{j}^{(i)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right)$

### v4

@@ 3.2 Bariant’s Attack
Recently Bariant introduced an attack on Ciminion via univariate greatest common divisor computations [Bar23]. As in our polynomial model, he introduced an auxiliary variable $X$ for the truncated component. If one inverts $p_{C}\circ p_{E}$ using $(c_{1}-p_{1},c_{2}-p_{2},X)^{\intercal}$ as output state, then one obtains three polynomials, see [Bar23, Fig. 4],
@@
Since the first polynomial is univariate in $X$, one can compute its greatest common divisor with $X^{q}-X$ to solve for $X$ and henceforth also solve for $K_{1}$ and $K_{2}$.
Observe that Bariant’s univariate polynomial is exactly the univariate polynomial of the LEX Gröbner basis.
Lemma 3.3.
Let $\mathbb{F}_{q}$ be a finite field, let $I_{\texttt{Ciminion}}\subset\mathbb{F}_{q}\big{[}y_{1},y_{2},\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(r-1)},x\big{]}$ be a Ciminion ideal, let $\tilde{f}\in\mathbb{F}_{q}[x]$ be the univariate LEX polynomial from Corollary 3.2, and let $f\in\mathbb{F}_{q}[x]$ be Bariant’s univariate polynomial. Then $f=\alpha\cdot\tilde{f}$ for some $\alpha\in\mathbb{F}_{q}^{\times}$.
Proof.
Ciminion is an iterated Feistel permutation therefore we have that
$\mathcal{R}^{(i)}\left(\mathbf{x}^{(i-1)}\right)\equiv\mathbf{x}^{(i)}\mod I_{\texttt{Ciminion}}\qquad\Longleftrightarrow\qquad\mathbf{x}^{(i-1)}\equiv{\mathcal{R}^{(i)}}^{-1}\left(\mathbf{x}^{(i)}\right)\mod I_{\texttt{Ciminion}},$
where $12\leq i\leq r-$ 1, and analogously for $i=1,r$. Since the inverse round functions are in $I_{\texttt{Ciminion}}$, we can iteratively substitute them into each other. This yields exactly the polynomials from Equation 32, so $f\in I_{\texttt{Ciminion}}$. By Corollary 3.2 we can construct a Ciminion LEX Gröbner basis in $x$-shape position, then for any $g\in\mathbb{F}_{q}[x]$ we have $g\in I_{\texttt{Ciminion}}\Leftrightarrow\tilde{f}\mid g$ [CLO15, Chapter 4 §5 Exercise 13]. Since $\deg\left(\tilde{f}\right)=\deg\left(f\right)$ they can only differ up to a multiplicative constant. ∎
Moreover, the construction of the univariate LEX polynomial can be effectively expressed in the cost of univariate polynomial multiplication over finite fields [CK91], which is quasi-linear in the degree.
As a countermeasure, Bariant proposes to add a linear combination of $K_{1}$ and $K_{2}$ to each component after $p_{C}$ [Bar23, §4]. We note that in general this modification breaks our Gröbner basis from Theorem 3.1. As a settlement, we propose to add a linear combination of $K_{1}$ and $K_{2}$ only to the non-linear parts of the round functions, i.e.
$\begin{split}\mathcal{R}^{(i)}:\mathbb{F}_{q}^{3}\times\mathbb{F}_{q}^{2}&\to\mathbb{F}_{q}^{3},\\ \begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}&\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&0&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+\alpha\cdot K_{1}+\beta\cdot K_{2}\end{pmatrix}+\mathbf{c}^{(i)},\end{split}$ (33)
where $\alpha,\beta\in\mathbb{F}_{q}^{\times}$. This modification does not affect the proof of Theorem 3.1, hence we maintain the DRL Gröbner basis for the modified Ciminion polynomial system. However, this modification breaks the construction of the univariate LEX polynomial via univariate polynomial multiplication, because the quadratic DRL Gröbner basis polynomial for the $r_{C}$th round will now also contain linear terms for the variables $x_{3}^{(2)},\dots,x_{3}^{(r-1)},x$ instead of only $x_{3}^{(r_{C}-1)},\dots,x_{3}^{(r-1)},x$. While the LEX Gröbner basis of the modified Ciminion might still be in $x$-shape position, we can now construct it only via an FGLM algorithm.
Note if the modification of Equation 33 is performed in all rounds of $p_{E}$, then the Middle State-Output Relation polynomial system of Bariant et al. [BBLP22, §5] becomes underdetermined, so this attack also becomes more difficult.
@@ 3.3 Implementation
We implemented Ciminion in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^5] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git The SageMath implementation can handle $128$ bit prime numbers, in particular the Gröbner basis from Theorem 3.1 can be computed for primes of that size. For $p=2^{127}+45$, $r_{C}=90$ and $r_{E}=14$ it is computed in negligible time.
@@ 3.3.1 Ciminion2
We also implemented a modified version of Ciminion to protect against Bariant’s attack, called Ciminion2, which adds $K_{1}+K_{2}$ in every non-linear round of $p_{E}$ and $p_{C}$, i.e. only in the first round of $p_{E}$ the key addition is omitted. The round function for the first sample then modifies to
$\mathcal{R}^{(i)}:\begin{pmatrix}x\\ y\\ z\end{pmatrix},\begin{pmatrix}K_{1}\\ K_{2}\end{pmatrix}\mapsto\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&1&1\end{pmatrix}\begin{pmatrix}x\\ y\\ z+x\cdot y+K_{1}+K_{2}\end{pmatrix}+\mathbf{c}^{(i)},$ (34)
@@
where $2\leq i\leq r_{C}+r_{E}$. For higher samples one then adds $K_{2\cdot j-1}+K_{2\cdot j}$ in $p_{E}$. For Ciminion2 a DRL Gröbner basis for the first sample is also computed in negligible time in SageMath for practical parameters. Due to the additional key addition the quadratic polynomials in the DRL Gröbner basis modify to
$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{3}^{(i)},\dots,x_{3}^{(r-1)},x\right)+\mathcal{A}_{i}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right),\qquad 2\leq i\leq r-1,$ (35)
$g_{3}^{(r)}$ $=\mathcal{Q}_{r}(x)+\mathcal{A}_{i}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right),$ (36)
where $\mathcal{Q}_{i}$ is quadratic and has the term ${x_{3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and $\mathcal{A}_{i}$ is affine and has non-zero coefficients for $x_{3}^{(2)},\dots,x_{3}^{(r-1)},x$.
3.4 Cryptanalysis
We now analyze the complexity of a Gröbner basis attack on Ciminion and Ciminion2 via our Gröbner basis, and compare it to the designer’s estimation and Bariant’s attack.
3.4.1 Estimation of the Designers
For Gröbner basis cryptanalysis of Ciminion [DGGK21, §4.4] the designers considered a weaker scheme where the keys are added after the permutations $p_{E}$ instead of before the rolling functions, see [DGGK21, Fig. 7]. The designers claimed that the weaker scheme is simpler to analyze than Ciminion. We reject this claim, after all we have computed a DRL Gröbner basis for Ciminion with rather simple linear variable eliminations, see Theorem 3.1.
The designers assumed that weakened Ciminion polynomial systems are regular, moreover for regular systems it is well-known that the degree of regularity $d_{\operatorname{reg}}$ is equal to the Macaulay bound (Theorem 2.6). Then, the DRL Gröbner basis complexity is bounded by [BFS04, Theorem 7]
$\mathcal{O}\left(\binom{n+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right)$ (37)
field operations. For Ciminion cryptanalysis, it is then conjectured that the DRL Gröbner basis computation is the most costly part in an attack.
The designers did not analyze the iterated model, they only studied a model of two equations in two key variables, we call this model the fully substituted model. By construction the polynomials in the fully substituted model have degree $2^{r_{C}+r_{E}-1}$. Under the regularity assumption we then have that $d_{\operatorname{reg}}=2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1=2^{r_{C}+r_{E}}-1$, so Equation 37 evaluates to
$\mathcal{O}\left(\binom{2^{r_{C}+r_{E}}+1}{2^{r_{C}+r_{E}}-1}^{\omega}\right)=\mathcal{O}\left(\left(\frac{2^{r_{C}+r_{E}}\cdot\left(2^{r_{C}+r_{E}}+1\right)}{2}\right)^{\omega}\right).$ (38)
In Table 3 this estimate is evaluated for sample we round numbers.
3.4.2 Complexity of Bariant’s Attack
Bariant’s attack is split into two parts: construction of the univariate polynomial via an iterated sequence of squarings and the GCD with the field equation. Multiplication of two univariate polynomials $f,g\in\mathbb{F}_{q}[x]$ with $\deg\left(f\right),\deg\left(g\right)\leq d$ can be performed in $\mathcal{O}\Big{(}d\cdot\log_{2}(d)\cdot\log_{2}\big{(}\log_{2}(d)\big{)}\Big{)}$ field operations [CK91], and the complexity for the GCD was given in Equation 21. The construction step then requires [Bar23, §3.2]
$\mathcal{O}\big{(}2^{r_{C}+r_{E}}\cdot(r_{C}+r_{E})\cdot\log_{2}(r_{C}+r_{E})\big{)}$ (39)
field operations, and the GCD requires
$\mathcal{O}\Big{(}2^{r_{C}+r_{E}-1}\cdot(r_{C}+r_{E}-1)\cdot\big{(}r_{C}+r_{E}-1+\log_{2}(q)\big{)}\cdot\log_{2}(r_{C}+r_{E})\Big{)}.$ (40)
Therefore, in general the GCD step dominates the complexity. In Table 3 Bariant’s attack is evaluated for sample we round numbers.
Note that the GCD step implicitly assumes that $2^{r_{C}+r_{E}-1}<q$, though if the degree of the Ciminion function exceeds $q$, then an exhaustive search over the truncated output always performs better than the GCD computation.
3.4.3 Eigenvalue Method
For standard Ciminion it is sufficient to find a $\mathbb{F}_{q}$-valued root for $x$, because we can then invert $p_{C}\circ p_{E}$ to obtain a key guess. Obviously, our Ciminion DRL Gröbner basis (Theorem 3.1) satisfies the structural assumption from Section 2.5.1, so the complexity of constructing a univariate polynomial in $x$ is, see Equation 20,
$\mathcal{O}\left(2^{\omega\cdot(r_{C}+r_{E}-2)+1}\right).$ (41)
However, for standard Ciminion this complexity is always dominated by Bariant’s attack, see Section 3.4.2.
For Ciminion2 on the other hand solving for $x$ is not sufficient, because we cannot invert $p_{C}\circ p_{E}$ anymore due to additional key additions. Therefore, for Ciminion2 we have to find a $\mathbb{F}_{q}$-valued solution for all quadratic variables to derive a key guess. Since the Ciminion2 DRL Gröbner basis satisfies the structure of Equation 22, the complexity of the Eigenvalue Method is given by Equation 25
$\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega+1}\cdot\frac{2^{\omega\cdot(r_{C}+r_{E}-1)}-N^{\frac{r_{C}+r_{E}-1}{2}}}{2^{2\cdot\omega}-N},&2\mid(r_{C}+r_{E}-1),\\ 2^{2\cdot\omega+1}\cdot\frac{2^{\omega\cdot(r_{C}+r_{E}-2)}-N^{\frac{r_{C}+r_{E}-2}{2}}}{2^{2\cdot\omega}-N},&2\nmid(r_{C}+r_{E}-1)\end{dcases}\end{rcases}\right),$ (42)
where $N$ is a bound on the number of $\mathbb{F}_{q}$-valued solutions in every iteration. In addition, we absorbed the terms for the GCDs with the field equations into the implied constant, because these steps are always dominated by the construction of the respective characteristic polynomial. Since a bound on $N$ is in general not known we consider a worst-case scenario approach. We assume that the first $\mathbb{F}_{q}$-valued solution that an adversary recovers in every iteration leads to the true key, i.e. we assume that $N=1$ in Equation 42.
In Table 3 we evaluated the discussed Ciminion complexity estimations. All our estimations are from a designer’s point of view, which assume an ideal adversary who has access to a hypothetical matrix multiplication algorithm which achieves $\omega=2$. As already mentioned, for standard Ciminion Bariant’s attack is the most performative attack, and at least $r_{C}+r_{E}\geq 112$ is required to achieve $128$ bits of security. For Ciminion2 on the other hand this attack is not feasible, so we have to fall back to the Eigenvalue Method. Then, we require at least $r_{C}+r_{E}\geq 66$ to achieve $128$ bits of security. Finally, the estimation of the Ciminion designers grossly underestimates the capabilities of Gröbner basis attacks. After all, we can construct an iterated Ciminion DRL Gröbner basis in negligible time, whereas the designers estimated this complexity to be in $\mathcal{O}\left(2^{2\cdot\omega\cdot(r_{C}+r_{E})}\right)$.
Complexity (bits)
$r_{C}+r_{E}$ Bariant’s Attack [Bar23] Eigenvalue Method Fully Substituted Model [DGGK21, §4.4]
$32$ $45.60$ $61.09$ $126$
$33$ $46.67$ $63.09$ $130$
$65$ $80.18$ $127.09$ $258$
$66$ $81.22$ $129.09$ $262$
$111$ $127.45$ $219.09$ $442$
$112$ $128.47$ $221.09$ $446$
Recall from [DGGK21, Table 1] that $r_{C}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$, where $s$ is the required security level. For Ciminion2 this already implies that $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion2}}\right)\geq 2^{\frac{s}{2}}$, and the complexity of the Eigenvalue Method will scale approximately with $\gtrsim 2^{\omega\cdot\frac{s}{2}}\geq 2^{s}$. So, for now we conclude that Ciminion2 resists Gröbner basis attacks within the security level, even under ideal assumptions on the adversary.
@@ 4 Analysis of Hydra
is linearly equivalent to the Feistel permutation
@@
via the matrices $\mathbf{A}_{1}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ and $\mathbf{A}_{2}=\begin{pmatrix}1&1\\ 0&1\\ \end{pmatrix}$, i.e. $\mathcal{L}=\mathbf{A}_{2}\circ\mathcal{F}\circ\mathbf{A}_{1}$. Inspired by this equivalence we can reduce the Hydra polynomial system to only one non-linear equation per round, i.e. the Hydra polynomial system transforms to a Feistel-like polynomial system.
In [Ste24, §6.2] the iterated polynomial systems representing generalized Feistel ciphers were analyzed. For these polynomial systems linear systems were derived, and in case the systems are of full rank the iterated polynomial systems are in generic coordinates, see [Ste24, Theorem 6.5]. For a system in generic coordinates we then have a proven upper bound on the solving degree, see Theorem 2.6.
Based on the aforementioned linear equivalence, we are going to replicate this strategy to construct a linear system for Hydra. Further, if the Hydra linear system has full rank, then we can transform the system into a DRL Gröbner basis via a linear change of coordinates.
@@ 4.1 Generic Coordinates for Hydra
Recall that the Hydra iterated polynomial system (Definition 2.2) consists of $816\cdot r_{\mathcal{H}}+$ 8 polynomials in $416\cdot r_{\mathcal{H}}+$ 4 variables. Let
$\mathbf{A}_{n}=\begin{pmatrix}1&0&\ldots&0&-1\\ 0&1&\ldots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&1&-1\\ 0&0&\ldots&0&1\end{pmatrix}\in\operatorname{GL}_{n}\left(\mathbb{F}_{q}\right),$ (45)
and let $\mathbf{B}=\operatorname{diag}\left(\mathbf{A}_{4},\mathbf{A}_{4}\right)\in\operatorname{GL}_{8}\left(\mathbb{F}_{q}\right)$. Then the first seven components of $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$, where $1\leq i\leq r_{\mathcal{H}}$ and $j\in\{1,2\}$, are affine and the eighth component is quadratic. Moreover, the fourth and the eighth components of $\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}$ are quadratic while the other ones are affine. In particular, by extracting the $2\cdot 7\cdot r_{\mathcal{H}}$ affine components of the $\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}$’s we can eliminate the variables
@@
where $10\leq i\leq r_{\mathcal{H}}-$ 1 and $j\in\{1,2\}$, in the quadratic equations. (This can be explicitly seen in the proof of Theorem 4.1.) Thus, after the elimination we reduce to $82\cdot r_{\mathcal{H}}+$ 8 polynomials, of which $6$ are also affine, in $42\cdot r_{\mathcal{H}}+$ 4 variables.
Next we derive a linear system to verify being in generic coordinates for Hydra.
@@
$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
@@
Then every homogeneous ideal $I\subset P[x_{0}]$ such that $x_{0}\notin\sqrt{I}$ and $\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\subset I$ is in generic coordinates if the following linear system has rank $416\cdot r_{\mathcal{H}}+$ 4
@@
$x_{j,k}^{(i-1)}-x_{j,8}^{(i-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i-1)}$ $=0,$
$x_{1,k}^{(r_{\mathcal{H}}-1)}-x_{1,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+(\mathbf{y},\mathbf{z})^{\intercal}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
$x_{2,k}^{(r_{\mathcal{H}}-1)}-x_{2,8}^{(r_{\mathcal{H}}-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{k}-\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}+\mathbf{x}_{2}^{(0)}\Big{)}\bigg{)}_{8}$ $=0,$
$\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}$ $=0,$
$y_{1}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{1}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}$ $=0,$
@@
$y_{3}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{3}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}$ $=0,$
@@
By [Ste24, Theorem 3.2] $\left(\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\right)$ is in generic coordinates if
$\sqrt{\mathcal{G}_{\textsf{Hydra}}^{\text{top}}}=\left(\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right).$ (47)